OpenStudy (anonymous):

Calculus - > Integral -> Substitution The question is sin 3x cos 3x dx Solve it by integrating using substitution.

OpenStudy (anonymous):

Do you know what the derivate of \[\Large \sin(3x)  \] is?

OpenStudy (anonymous):

derivative*

geerky42 (geerky42):

HINT: Let u be sin(3x).

OpenStudy (anonymous):

How do you choose what to let u be? In this case you choose sin(3x) but why?

OpenStudy (anonymous):

Takes some practice, but in case you consider taking u=cos(3x), that would also work. It's mainly about developing the idea of substitution. If you choose u=sin(3x) then: \[\Large \frac{du}{dx}=3\cos(3x) \]